Fowler and Woods Enterprises is a publicly traded company that just paid a $2.00 per share dividend.  The company is expected to increase its dividend by 20% per year for the next two years.  After the second year, the dividend growth rate will be 5% per year for the next two years.  After the 4th year, dividends are expected to grow at a constant rate of 3% into the foreseeable future.  An analyst estimates that investors in the firm will require a 12% annual return.  Based on this information, what is the intrinsic value of the stock today?
Answer #1

Given, D0=2

D1=D0*(1+20%)=2*1.2=2.4

D2=D1*(1+20%)=2.4*1.2=2.88

D3=D2*(1+5%)=2.88*1.05=3.024

D4=D3*(1+5%)=3.024*1.05=3.175

D5=D4*(1+3%)=3.175*1.03=3.27

Horizon value at year4=D5/(required rate-growth rate)

=3.27/(12%-3%)

=36.338

Value of the stock=(D1/(1+12%))+(D2/(1+12%)^2)+(D3/(1+12%)^3)+((D4+Horizon value)/(1+12%)^4)

=(2.4/1.12)+(2.88/1.12^2)+(3.024/1.12^3)+(39.51/1.12^4)

=2.14+2.30+2.15+25.11

=$31.70

Value of the stock=$31.70